OpenStudy (anonymous):

Solve the equation for x. 3x = –4y

OpenStudy (anonymous):

Isolate x. 3x = -4y Divide both sides by 3. x = -4/3y

OpenStudy (anonymous):

We divide by 3 because... 3 and x are being multiplied: 3x = 3 * x We want to find x alone, so we have to undo 3x... 3(x) /x = 3